OpenStudy (anonymous):

Please help me out?:) Find u x v and show that it is orthogonal to both u and v. u=6k, v=-i+3j+k . thx

OpenStudy (anonymous):

I think the answer is (-18,-6,0) but idk how to prove that it is orthogonal

OpenStudy (kainui):

Well if you're taking the cross product...

OpenStudy (fwizbang):

take your answers dot product with u and with v. If you get zero, its orthogonal(perpendicular)

OpenStudy (blockcolder):

Compute the dot product of u X v and u, and then u X v and v. Both of them should be 0.

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

and is my answer right?

OpenStudy (blockcolder):

\[(\mathbf{u} \times \mathbf{v})\cdot \mathbf{u}=(-18)(0)+(-6)(0)+(0)(6)=0\\ (\mathbf{u} \times \mathbf{v}) \cdot \mathbf{v}=(-18)(-1)+(-6)(3)+(0)(1)=0\] Thus, u X v is orthogonal to both u and v.

OpenStudy (anonymous):

ok thanks. is my answer correct

OpenStudy (phi):

yes, your cross product is correct

OpenStudy (anonymous):

thx